what is the mass of 4.9 * 10^26 molecules of sodium chloride?


1
Expert's answer
2021-01-11T03:46:55-0500

One mole of sodium chloride molecules i.e. 6.022 * 1023 molecules of NaCl weigh 58.44g.

no. of moles in 4.9 * 1026 molecules ="\frac{4.9*1000}{6.022}" = 813.683 moles.

Mass of 813.638 moles = 47551.644g